OpenStudy (anonymous):

The backdrop for a music concert is rectangular in shape. A satin ribbon of length 30 feet is draped along the diagonal of the backdrop. The length of the backdrop is 24 feet. Which expression can be used to find the height of the backdrop?

OpenStudy (ashleyisakitty):

-a^2+b^2=c^2. -24^2+b^2=30^2 -576+b^2=900 -900-576=b^2 -324=b^2 then you square root 234 sooo an expression would be sqrt of 30^2 - 24^2.